Solve linear equation.check.
3x/4+6=-12
Thank you for any help I receive.

Answer by scott8148(6628)

subtracting 6 ___ 3x / 4 = -18

dividing by 3 ___ x / 4 = -6

multiplying by 4 ___ x = -24

checking ___ 3(-24) / 4 + 6 = -12

-72 / 4 + 6 = -12

-18 + 6 = -12

-12 = -12